Two charges each with a magnitude of1.1

â‹…

1


−

7

1.1⋅10−7

C experience a force of 4.2

â‹…

1


−

4

4.2⋅10−4

N. How far apart are the charges from each other? 


1
Expert's answer
2021-05-20T18:24:16-0400

According to the Coulomb's law, the force experienced by each charge is:


"F = k\dfrac{q^2}{r^2}"

where "q = 1.1\times 10^{-7}C" is the charge of both charges, "r" is the distance between charges, and "k = 9\times10^{9}\space N\cdot m^2/C^2" is the Coulomb's constant. Expressing "r" and substituting "F = 4.2\times 10^{-4}N", obtain:


"r = q\sqrt{\dfrac{k}{F}}\\
r = 1.1\times 10^{-7}\cdot \sqrt{\dfrac{9\times10^{9}}{4.2\times 10^{-4}}} \approx 0.51m"

Answer. 0.51 m.